# TempoControl: Temporal Attention Guidance for Text-to-Video Models

### v1

@@ Abstract
Recent advances in generative video models have enabled the creation of high-quality videos based on natural language prompts. However, these models frequently lack fine-grained temporal control, meaning they do not allow users to specify when particular visual elements should appear within a generated sequence. In this work, we introduce TempoControl, a method that allows for temporal alignment of visual concepts during inference, without requiring retraining or additional supervision. TempoControl utilizes cross-attention maps, a key component of text-to-video diffusion models, to guide the timing of concepts through a novel optimization approach. Our method steers attention using three complementary principles: aligning its temporal shape with a control signal (via correlation), amplifying it where visibility is required (via energy), and maintaining spatial focus (via entropy). TempoControl allows precise control over timing while ensuring high video quality and diversity. We demonstrate its effectiveness across various video generation applications, including temporal reordering for single and multiple objects, as well as action and audio-aligned generation.
@@
Codehttps://github.com/Shira-Schiber/TempoControl
Project Pagehttps://shira-schiber.github.io/TempoControl/
@@ 1 Introduction
Generative video models have improved dramatically, enabling the synthesis of high-quality, temporally coherent videos from natural language descriptions (Singer et al. 2022; Blattmann et al. 2023; Yang et al. 2024; Bar-Tal et al. 2024; Wan et al. 2025). These models have enabled creative applications in animation, design, and virtual content creation by generating visually compelling sequences that align with textual descriptions. However, while current models offer strong spatial consistency and global semantics, they lack fine-grained temporal control. Specifically, guiding the model to generate visual elements at specific time points is particularly challenging. This includes making an object appear in the middle of a scene or aligning visual aspects of lighting with the sound of thunder (see Figure 1).
Recent works on generative video modeling have proposed various spatial and motion control mechanisms, including object masking, layout conditioning, or added camera controls (Gupta et al. 2024; Wang et al. 2024b; Feng et al. 2025; He et al. 2024). In contrast, temporal control has received significantly less attention. Incorporating such control typically requires additional supervision in the form of temporally annotated datasets. Since video data is often scarce, extending annotations to capture temporal variation is expensive and often impractical at scale. Synthetic data could be used; however, generating diverse and temporally accurate video-text pairs, especially for abstract concepts such as movement, is a highly non-trivial challenge.
In this work, we propose a lightweight yet effective inference-time control method, TempoControl, which introduces temporal control for generated video. Instead of retraining the generative model, TempoControl leverages a property already embedded in these systems: the text-frame cross-attention mechanism.
@@
Attention control is achieved by carefully steering the latent variables during the diffusion process. At each denoising step, we apply a few stochastic gradient descent iterations (Battash, Wolf, and Lindenbaum 2024) until a satisfactory level of temporal alignment is achieved, notably, without updating model parameters.
@@
To guide temporal alignment, TempoControl introduces a novel loss with three complementary components, based on: Pearson correlation, energy, and entropy. The Pearson term encourages the temporal shape of attention to match that of an external control signal, operating on the normalized attention vector to align the presence of a concept with the desired timing. However, since correlation is scale-invariant, it may yield high scores even when attention values are too low to render the object visible. To mitigate this, we incorporate an energy term that directly promotes stronger attention in frames where the temporal signal is high, and suppresses it elsewhere. Additionally, we introduce an entropy regularization term to maintain spatial focus, ensuring that when attention is activated, it remains coherent and not diffusely spread across the frame.
@@
Beyond single-object scenarios, as shown in Figure 1, our method generalizes well to multi-object scenes. In such cases, it is important to maintain the visual grounding of non-target entities; we find that preserving one object (e.g., the bird) while selectively manipulating another (e.g., the cat) helps retain the overall scene structure. In addition to nouns, our method can also manipulate verbs (e.g., clapping), effectively modifying the motion depicted in the video. Finally, we explore the potential of aligning external audio cues with video generation. While we focus on simple cases involving a single audio event (e.g., a lightning strike), we observe promising results, suggesting that our approach may naturally extend to zero-shot multimodal control without additional training.
@@
Our contributions are as follows: (i) We propose an inference-time method for temporally controlling the appearance of specific words in text-to-video diffusion models, without requiring model finetuning or additional supervision. (ii) We introduce a suite of losses that operate over spatiotemporal attention to enable fine-grained control over both the timing and intensity of word grounding. (iii) We demonstrate the versatility of our method across various settings, including multi-object control, motion manipulation, and audio-to-video alignment, and introduce a new metric to evaluate the task of temporal alignment control, which is not covered by existing benchmark suites.
Figure 2: Video generated for the prompt “A cat and a dog.” Before (left) and after optimization (right). Top: Spatial attention maps for the tokens cat and dog at denoising step t=3t{=}3, shown for frames j=2,10,19j{=}2,10,19. Attention is overlaid on generated frames. Bottom: temporal attention $a_{i,j}^{t}$ (blue) vs. target mask $m_{i,j}$ (orange), with Pearson correlation loss shown.
@@ 2 Related Work
Motion Control. Fine-grained motion control has been an active area of research. Boximator (Wang et al. 2024b) introduces control by allowing users to specify the spatial locations of a start box and an end box for each object. This control is implemented by fine-tuning a dedicated control module. BlobGEN-Vid (Feng et al. 2025) enhances control by introducing visual primitive-based guidance, such as lines, to enable detailed manipulation of object motion, fine-grained appearance, and smooth object transitions. CameraCtrl (He et al. 2024) addresses controllability by enabling precise camera movement and perspective control. This is accomplished by training a plug-and-play camera pose control module on top of a video diffusion model. Motion Prompting (Geng et al. 2025) enables movement control by incorporating spatio-temporally sparse or dense motion trajectory inputs. However, these approaches require extensive datasets and computational resources for training, substantial user annotations during inference, and notably lack precise control over motion timing or peak movement appearance.
Attention-based Control. Attention-based methods have improved spatial control in text-to-image generation. Prompt-to-Prompt (Hertz et al. 2023) enables word-level edits via cross-attention manipulation, while Attend-and-Excite (Chefer et al. 2023) ensures all mentioned objects are depicted by balancing token attention. Battash et al. (2024) introduced a layout optimization strategy for multi-subject generation in text-to-image diffusion, which constrains cross-attention to improve the spatial separation of objects. Be Yourself (Dahary et al. 2024) uses bounded attention to preserve subject identity and prevent semantic leakage in multi-subject text-to-image generation. Stable Flow (Avrahami et al. 2025) enables training-free image editing by selectively injecting attention features into vital layers of a diffusion transformer. These methods operate on static images, whereas we extend attention-based control to the temporal domain, enabling control over object timing and motion in video generation.
@@
Inference time optimization. Inference-time optimization for text-to-video diffusion has only recently attracted attention. (Li et al. 2025) pre-compute a spatio-temporal sketch and inject plan-consistent noise during sampling to tighten prompt alignment. (Wei et al. 2025) converts a single keyframe into a textured, deformable 3D mesh that is re-rendered at every denoising step to maintain consistency in the garment’s appearance. (Shaulov et al. 2025) derive variance-based optical-flow cues from the model’s latent predictions and add a flow-coherence loss on-the-fly, smoothing motion without retraining. (Lian et al. 2023) let a large language model generate scene graphs and temporal layouts that act as attention masks, improving complex spatio-temporal grounding. While these methods improve fidelity and consistency without necessitating additional training, none offer explicit, user-defined control over the timing of textual element unfolding.
@@ 3 Preliminaries
We build TempoControl on top of Wan 2.1 (Wan et al. 2025), a state-of-the-art diffusion model that serves as our backbone for text-to-video generation. Let $v\in\mathbb{R}^{T_{v}\times H_{v}\times W_{v}\times 3}$ represent the input video, where $T_{v}$ is the number of frames, $H_{v}\times W_{v}$ denotes the spatial resolution of each frame, and the final dimension corresponds to the three RGB color channels. The video is first encoded by a 3D causal variational autoencoder into a latent tensor
@@
3.1 Cross-attention in text-to-video diffusion models
A key challenge in text-to-video generation is ensuring that the synthesized video aligns semantically and temporally with the input prompt. In diffusion-based architectures, this alignment is achieved through cross-attention layers within the transformer blocks at each denoising step $t$.
@@ 3.1 Cross-attention in text-to-video diffusion models
The effect of a specific word $i$ is measured by extracting the corresponding column $\bar{A}^{t}_{i}\in\mathbb{R}^{n_{v}}$. The vector $\bar{A}^{t}_{i}$ can be reshaped at temporal index $j$, yielding $\bar{A}^{t}_{j,i}\in\mathbb{R}^{H^{\prime}\times W^{\prime}}$, which represents the spatial influence of word $i$ at temporal position $j$ and denoising step $t$. Figure 2 illustrates this for the words cat and dog in the prompt “A cat and a dog.” The attention maps are up-sampled to the frame resolution for visualization.
The attention map captures the spatiotemporal grounding of words in the generated video; TempoControl is specifically designed to steer the temporal influence of a token. Therefore, we focus on the scalar attention value $\hat{A}^{t}_{j,i}=\langle\bar{A}^{t}_{j,i}\rangle_{x,y}\in\mathbb{R}$, where $\langle\cdot\rangle_{x,y}$ denotes summation over the spatial coordinates $x\in[1,H^{\prime}]$ and $y\in[1,W^{\prime}]$. Intuitively, $\hat{A}^{t}_{j,i}$ measures the influence of word $i$ on the latent representation at temporal position $j$ during the denoising step $t$.
To summarize the temporal attention pattern of word $i$ at denoising step $t$, we collect the scalar values across frames into a vector:
@@
We omit the first frame due to unstable attention values that do not reliably reflect word grounding. This vector serves as the target for our steering method, described in the next section.
@@ 4 Method
Given a video $v$, our goal is to generate an edited version $v^{\prime}$ such that specific tokens appear according to a temporal condition defined by a weighting mask $m_{i}=[m_{i,1},\ldots,m_{i,T^{\prime}}]$, where $m_{i,j}\in[0,1]$ indicates the desired strength of activation of token $p_{i}$ at temporal position $j$ (see Figure 2 bottom). In most experiments, we use binary masks (i.e., values of zero or one), except for the audio-alignment setting, where the mask reflects sound strength. The main steps of our approach are detailed in the appendix Algorithm 1, illustrated in Figure 3, and discussed below.
We steer the latent representation during the first $k$ denoising steps by performing $l$ gradient update iterations at each step using the AdamW optimizer. For clarity, the update can be expressed as:
@@
where $\alpha\in\mathbb{R}$ is the learning rate and $t$ is the current denoising step. The loss $\mathcal{L}^{t}$ consists of four components (detailed below): (i) temporal correlation, (ii) energy, (iii) entropy regularization, and (iv) spatial consistency.
@@
The first loss term encourages the temporal attention vector $a_{i}^{t}$ to match the target mask $m_{i}$ using the Pearson correlation:
@@
where $\tilde{a}_{i}^{t}$ is the min-max normalized version of $a_{i}^{t}$, defined as $\tilde{a}_{i}^{t}=\frac{a_{i}^{t}-\min(a_{i}^{t})}{\max(a_{i}^{t})-\min(a_{i}^{t})}$. Here, $\text{Cov}(\cdot,\cdot)$ denotes empirical covariance, and $\sigma_{m_{i}}$ and $\sigma_{\tilde{a}_{i}^{t}}$ are the corresponding standard deviations. Note that if $m_{i}$ has zero variance, such as when encouraging an object to appear throughout the entire video, the correlation term is undefined; in such cases, it is excluded from the loss.
This alignment term adjusts the temporal placement of words in the video but does not affect their overall strength. As a result, words that are initially invisible may remain so. The following loss term is designed to modulate the temporal strength of attention.
Attention energy term.
The energy term adjusts the overall magnitude of attention for word token $i$ based on the temporal mask $m_{i}$. We use the indicator function $\mathbb{1}_{\{\cdot\}}$, which equals 1 when the condition is true and 0 otherwise. The positive energy encourages attention where the word should be active:
@@
The negative energy penalizes attention where the word should not be visible:
@@
The final energy loss is defined as the net difference:
$\mathcal{L}_{\text{energy}}^{t}=\mathcal{L}_{\ominus}^{t}-\mathcal{L}_{\oplus}^{t}.$ (7)
Here, $\tau\in[0,1]$ is a threshold that determines whether a given time index is considered active for token $p_{i}$. While $m_{i}$ is often binary (i.e., 0 or 1), it may also take continuous values (e.g., for soft audio-driven conditioning). In such cases, the indicator $\mathbb{1}_{\{m_{i,j}>\tau\}}$ selects the time-steps where the target activation is sufficiently strong.
Increasing the attention magnitude without constraint can lead to overly diffuse spatial maps (Qi et al. 2025), especially in multi-object scenes, which degrades the visual grounding of word tokens. To regularize this behavior and preserve spatial consistency, we introduce additional loss terms in the following paragraph.
Entropy regularization.
Word-level attention is typically spatially focused rather than dispersed (see Figure 2). To discourage excessive spreading of attention during optimization, we regularize the entropy of the spatial attention maps. Specifically, we apply this regularization only at temporal indices where the word is expected to be active (i.e., $m_{i,j}>\tau$), resulting in:
@@
Spatial consistency penalty.
To maintain spatial consistency, we penalize deviations between the initial and optimized spatial attention maps for the target word. Here, $\hat{A}^{t}_{j,i}$ denotes the spatial attention map before optimization at step $t$, for word $i$ at frame $j$. The penalty is defined as
$\mathcal{L}_{\text{spatial}}^{t}=\left\|\sum_{j=1}^{T^{\prime}}\left(\hat{A}^{t}_{j,i}-\bar{A}^{t}_{j,i}\right)\right\|_{2}^{2},$ (9)
where the difference is computed over the spatial dimensions, after summing across all latent frames.
@@
$\mathcal{L}^{t}=\mathcal{L}_{\text{corr}}^{t}+\lambda_{1}\mathcal{L}_{\text{energy}}^{t}+\lambda_{2}\mathcal{L}_{\text{entropy}}^{t}+\lambda_{3}\mathcal{L}_{\text{spatial}}^{t},$ (10)
where $\lambda_{1},\lambda_{2},\lambda_{3}$ are hyperparameters used to calibrate the contribution of the regularization terms. This objective is optimized over latent codes during inference; model weights remain fixed.
@@
Method Temporal Accuracy T. Absence Accuracy T. Presence Accuracy Subject Consistency Background Consistency Motion Smoothness Dynamic Degree Aesthetic Quality Imaging Quality One Object Text 63.94% 67.38% 60.50% 84.66% 92.30% 99.16% 47.50% 45.98% 53.76% Ours 81.00% 88.75% 72.91% 80.60% 90.33% 99.05% 38.75% 44.37% 52.99% Two Objects Text 37.50% 45.85% 29.15% 92.06% 95.32% 98.92% 62.20% 53.26% 68.56% Ours 55.00% 66.46% 43.54% 92.21% 95.43% 99.19% 41.46% 56.55% 70.66% Movement Text 19% 93.17% 95.91% 98.10% 73.00% 55.81% 60.46% Ours 53% 92.86% 95.67% 98.20% 80.00% 55.06% 61.76%
Table 1: Comparison against Wan 2.1 across three temporal-control applications: (i) one object, (ii) two objects, and (iii) movement. Results for irrelevant metrics are marked with –. Higher values indicate better performance for all metrics.
Method Multiple Object (GriT) Multiple Object (YOLO) Dynamic Degree Text 74.13% 61.54% 30.49% Ours 76.37% 65.73% 18.78%
Table 2: Multiple object benchmark assessing the ability to depict multiple objects across the video.
Question Type Ours Baseline Same Temporal Accuracy 61.51% 16.94% 21.55% Visual Quality 62.66% 25.99% 11.35%
Table 3: User study results for temporal accuracy and visual quality.

### v2

@@ Abstract
Recent advances in generative video models have enabled the creation of high-quality videos based on natural language prompts. However, these models frequently lack fine-grained temporal control, meaning they do not allow users to specify when particular visual elements should appear within a generated sequence. In this work, we introduce TempoControl, a method that allows for temporal alignment of visual concepts during inference, without requiring retraining or additional supervision. TempoControl utilizes cross-attention maps, a key component of text-to-video diffusion models, to guide the timing of concepts through a novel optimization approach. Our method steers attention using three complementary principles: aligning its temporal pattern with a control signal (correlation), adjusting its strength where visibility is required (magnitude), and preserving semantic consistency (entropy). TempoControl provides precise temporal control while maintaining high video quality and diversity. We demonstrate its effectiveness across various applications, including temporal reordering of single and multiple objects, action timing, and audio-aligned video generation. Please see our project page for more details: https://shira-schiber.github.io/TempoControl/.
@@ 1 Introduction
Generative video models have improved dramatically, enabling the synthesis of high-quality, temporally coherent videos from natural language descriptions [22, 6, 34, 3, 23]. These models have enabled creative applications in animation, design, and virtual content creation by generating visually compelling sequences that align with textual descriptions. However, while current models offer strong spatial consistency and global semantics, they lack fine-grained temporal control. Specifically, guiding the model to generate visual elements at specific time points is particularly challenging. This includes making an object appear in the middle of a scene or aligning visual aspects of lighting with the sound of thunder (see Figure 1).
Recent works on generative video modeling have proposed various spatial and motion control mechanisms, including object masking, layout conditioning, or added camera controls [10, 25, 9, 11]. In contrast, temporal control has received significantly less attention. Incorporating such control typically requires additional supervision in the form of temporally annotated datasets. Since video data is often scarce, extending annotations to capture temporal variation is expensive and usually impractical at scale. Synthetic data could be used; however, generating diverse and temporally accurate video-text pairs, especially for abstract concepts such as movement, is a highly non-trivial challenge.
In this work, we propose a lightweight yet effective inference-time control method, TempoControl, which introduces temporal control for generated video. Instead of retraining the generative model, TempoControl leverages a property already embedded in these systems: the text-vision cross-attention mechanism.
@@
Attention control is achieved by carefully steering the latent variables during the diffusion process. At each denoising step, we apply a few stochastic gradient descent iterations [7, 5] until a satisfactory level of temporal alignment is achieved, notably, without updating model parameters.
@@
To guide temporal alignment, TempoControl introduces a novel loss with three complementary components, based on Pearson correlation, magnitude, and entropy. The Pearson term encourages the temporal shape of attention to match that of an external control signal, operating on the normalized attention vector to align the presence of a concept with the desired timing. However, since correlation is scale-invariant, it may yield high scores even when attention values are too low to render the object visible. To mitigate this, we incorporate a magnitude term that directly promotes stronger attention in frames where the temporal signal is high, and suppresses it elsewhere. Additionally, we introduce an entropy regularization term to maintain spatial focus, ensuring that when attention is activated, it remains coherent and not diffusely spread across the frame.
@@
Beyond single-object scenarios, as shown in Figure 1, our method generalizes well to multi-object scenes. In such cases, it is crucial to maintain the visual grounding of non-target entities. We find that preserving one object (e.g., the bird) while selectively manipulating another (e.g., the cat) helps retain the overall scene structure. In addition to nouns, our method can also manipulate verbs (e.g., clapping), effectively modifying the motion depicted in the video. Finally, we explore the potential of aligning external audio cues with video generation. While we focus on simple cases involving a single audio event (e.g., a lightning strike), we observe promising results, suggesting that our approach may naturally extend to zero-shot multimodal control without additional training.
@@
Our contributions are threefold: (i) We introduce the first inference-time method for temporally controlling the appearance of specific words in text-to-video diffusion models, without additional supervision or model finetuning, allowing temporal control while preserving the fidelity and diversity of existing models. (ii) We propose a set of spatiotemporal attention losses that shape both the timing and strength of word grounding, together with an entropy-based regularization term that maintains semantic consistency. (iii) We demonstrate broad applicability across multi-object control, motion timing, and audio-to-video alignment, and introduce a new metric for temporal alignment, a dimension not covered by current benchmarks.
Figure 2: Motivation for our approach. Top: We show a video generated for the prompt “The video begins with a serene view centered on the cat, with no sign of the dog. In the second half, the dog unexpectedly appears, altering the dynamic of the scene.” The top row displays the attention maps for the tokens cat and dog, extracted from the denoising step t=3t{=}3, for frames j=2,6,10,14,19j{=}2,6,10,14,19. On the left, the video is generated without our optimization. Despite the prompt specifying that the dog should appear in the second half, it appears early. This behavior is common, as Wan 2.1 often fails to depict objects or movements according to temporal cues in the prompt. On the right, after applying our conditioning method, the dog correctly appears in the second half of the video. Bottom: Temporal attention $a_{i,j}^{t}$ (blue) versus target mask $m_{i,j}$ (orange), with the corresponding Pearson correlation loss shown.
@@ 2 Related Work
Attention-based Control. Attention manipulation has become a powerful tool for controlling semantics in text-to-image generation. Methods such as Attend-and-Excite [7, 17, 4, 8] perform inference-time optimization over cross-attention maps to ensure that all textual entities are faithfully represented in the generated image. Stable Flow [1] further introduces a training-free editing framework that selectively injects attention features into key layers of a diffusion transformer. Additional approaches, including Prompt-to-Prompt [12], leverage attention restructuring for fine-grained, text-driven edits.
@@
Inference time optimization. Inference-time optimization for text-to-video diffusion has only recently attracted attention. [16] pre-compute a spatio-temporal sketch and inject plan-consistent noise during sampling to tighten prompt alignment. [28] converts a single keyframe into a textured, deformable 3D mesh that is re-rendered at every denoising step to maintain consistency in the garment’s appearance. [21] derive variance-based optical-flow cues from the model’s latent predictions and add a flow-coherence loss on-the-fly, smoothing motion without retraining. [18] let a large language model generate scene graphs and temporal layouts that act as attention masks, improving complex spatio-temporal grounding. While these methods improve fidelity and consistency without necessitating additional training, none offer explicit, user-defined control over the timing of textual element unfolding.
Controllable Video Generation Controllable video generation has advanced rapidly, with methods exploring a broad set of control modalities. Early multimodal systems, such as VideoComposer [26], incorporate depth, sketch, and motion vector guidance to steer video synthesis. More recent approaches introduce explicit camera and object motion control on top of text-to-video generation [27, 11, 15, 33, 9, 2]. Another line of work investigates the controllability of image-to-video object motion [25, 32]. While effective, these training approaches rely on large paired datasets and extensive supervision, which can be costly and limit generalization.
A parallel line of work studies training-free controllability, including CamTrol [13] for camera motion, MotionClone and DiTFlow [19, 20] for transferring motion from reference videos, and MOFT [31] for editing motion at inference. These rely on dense motion trajectories, but cannot specify when a concept should appear or disappear.
The most related work, MinT [30], arranges multiple textual instructions in temporal order by learning when each should begin. However, MinT provides only coarse event sequencing, whereas our method offers fine-grained timing for individual concepts. Moreover, MinT requires fine-tuning on temporally annotated caption datasets centered on human activities, which limits its generality across diverse prompts. Our approach operates entirely at inference time and generalizes across settings without any training or domain-specific supervision.
@@ 3 Preliminaries
We build TempoControl on top of Wan 2.1 [23], a state-of-the-art diffusion model that serves as our backbone for text-to-video generation. Let $v\in\mathbb{R}^{T_{v}\times H_{v}\times W_{v}\times 3}$ represent the input video, where $T_{v}$ is the number of frames, $H_{v}\times W_{v}$ denotes the spatial resolution of each frame, and the final dimension corresponds to the three RGB color channels. A 3D causal variational autoencoder first encodes the video into a latent tensor
@@
3.1 Cross-attention in video diffusion models
A key challenge in text-to-video generation is maintaining semantic and temporal alignment with the input prompt. In diffusion-based architectures, this alignment is achieved through cross-attention layers within the transformer blocks at each denoising step $t$.
@@ 3.1 Cross-attention in video diffusion models
The effect of a specific word $i$ is measured by extracting the corresponding column $\bar{A}^{t}_{i}\in\mathbb{R}^{n_{v}}$. The vector $\bar{A}^{t}_{i}$ can be reshaped at temporal index $j$, yielding $\bar{A}^{t}_{j,i}\in\mathbb{R}^{H^{\prime}\times W^{\prime}}$, which represents the spatial influence of word $i$ at temporal position $j$ and denoising step $t$. Figure 2 illustrates this for the words cat and dog in the prompt “A cat and a dog. In the second half of the video, the dog appears.” Attention maps are upsampled to frame resolution for display.
The attention map captures the spatiotemporal grounding of words in the generated video; TempoControl is specifically designed to steer the temporal influence of a token. Thus, we focus on the scalar attention value $\hat{A}^{t}_{j,i}=\langle\bar{A}^{t}_{j,i}\rangle_{x,y}\in\mathbb{R}$, where $\langle\cdot\rangle_{x,y}$ denotes summation over the spatial coordinates $x\in[1,H^{\prime}]$ and $y\in[1,W^{\prime}]$. Intuitively, $\hat{A}^{t}_{j,i}$ measures the influence of word $i$ on the latent representation at temporal position $j$ during denoising step $t$.
To aggregate the temporal attention pattern of word $i$ at denoising step $t$, we collect the scalar values across frames into a vector:
@@
We omit the first frame due to unstable attention values that do not reliably reflect word grounding. This vector serves as the target for our steering method, described next.
@@ 4 Method
Given a video $v$, our objective is to generate an edited version $v^{\prime}$ such that specific tokens appear according to a temporal condition. To achieve fine-grained temporal control, we employ masks, which provide a more precise and less ambiguous mechanism than textual descriptions. We denote as input a binary mask $m_{i}=[m_{i,1},\ldots,m_{i,T^{\prime}}]$, where each element $m_{i,j}\in\{0,1\}$ indicates whether token $p_{i}$ should appear at temporal position $j$ (see Figure 2, bottom, orange line). In this example, the mask specifies that the token dog should appear only in the second half of the video (i.e., frames 10–20). In general, $m_{i,j}$ may also take continuous values to represent the strength of appearance. In most experiments, we use binary masks, except in the audio-alignment setting, where the mask encodes sound intensity. The main steps of our approach are illustrated in Figure 3 and discussed below. We also provide an algorithmic description in the appendix.
We steer the latent representation during the first $k$ denoising steps by applying $l$ gradient-based updates at each step, optimized with AdamW. Let $z_{t}$ denote the latent at step $t$; the update is:
@@
where $\alpha$ is the learning rate. The loss $\mathcal{L}^{t}$ consists of a single main component: a temporal correlation term that aligns the attention pattern with the target mask. While this term is most significant for aligning with the temporal condition, we also include two terms that further improve the results: (ii) a magnitude term that amplifies or suppresses token-level activation, and (iii) an attention entropy regularization term that preserves spatial consistency. We describe each component in detail below.
@@
The main loss term encourages the temporal attention vector $a_{i}^{t}$ to match the target mask $m_{i}$ using the Pearson correlation:
@@
where $\tilde{a}_{i}^{t}$ is the min–max normalized version of $a_{i}^{t}$, defined as $\tilde{a}_{i}^{t}=\frac{a_{i}^{t}-\min(a_{i}^{t})}{\max(a_{i}^{t})-\min(a_{i}^{t})}$. Here, $\text{Cov}(\cdot,\cdot)$ denotes the empirical covariance, and $\sigma_{m_{i}}$ and $\sigma_{\tilde{a}_{i}^{t}}$ are the corresponding standard deviations. Note that if $m_{i}$ has zero variance, such as when encouraging an object to appear throughout the entire video, the correlation term becomes undefined; in such cases, it is omitted from the loss.
The Pearson correlation measures the linear relationship between the mask $m_{i}$ and the attention term $a_{i}^{t}$, but it does not capture their scale. In our experiments, we observe two phenomena: (i) words that are initially invisible may remain so, and (ii) completely suppressing the appearance of a token is difficult. The following loss term is therefore designed to modulate the temporal strength of attention.
Attention magnitude term.
The magnitude term adjusts the overall strength of attention for each word token $i$ according to the temporal mask $m_{i}$, complementing the correlation-based term by restoring scale sensitivity. We employ the indicator function $\mathbb{1}_{\{\cdot\}}$, which equals one when the specified condition holds and zero otherwise. The positive magnitude component encourages attention in regions where the word is expected to be active:
@@
The negative magnitude component penalizes attention where the word should not be visible:
@@
The final magnitude loss is defined as the net difference:
$\mathcal{L}_{\text{mag}}^{t}=\mathcal{L}_{\ominus}^{t}-\mathcal{L}_{\oplus}^{t}.$ (7)
Figure 4: Entropy regularization helps preserve semantic fidelity.
Here, $\tau\in[0,1]$ is a threshold that determines whether a given time index is considered active for token $p_{i}$. While $m_{i}$ is often binary (i.e., 0 or 1), it may also take continuous values (e.g., for soft audio-driven conditioning). In such cases, the indicator $\mathbb{1}_{\{m_{i,j}>\tau\}}$ selects time steps where the target activation is sufficiently strong.
Both the correlation and magnitude terms target the temporal objective but can harm spatial consistency. In our experiments, these terms sometimes corrupt object semantics, such as a smartphone turning into a distorted two-screen tablet (Figure 4; more examples in the appendix). To address this and preserve spatial coherence, we introduce an entropy-based regularization term.
Attention entropy regularization.
Optimizing for the temporal objective may compromise spatial properties, often resulting in overly spread attention. We counter this by regularizing the entropy of the spatial attention maps at temporal indices where the word is active (i.e., $m_{i,j}>\tau$), resulting in:
@@
In our experiments, we find that discouraging excessive spreading of attention not only improves temporal presence but also enhances overall image quality (see ablations), suggesting that this term may be beneficial for video generation more broadly.
@@
$\mathcal{L}^{t}=\mathcal{L}_{\text{corr}}^{t}+\lambda_{1}\mathcal{L}_{\text{magnitude}}^{t}+\lambda_{2}\mathcal{L}_{\text{entropy}}^{t},$ (9)
where $\lambda_{1}$ and $\lambda_{2}$ control the auxiliary terms. The correlation term is sufficient for most cases and provides strong temporal alignment. The magnitude term is helpful when the base model fails to accurately depict a token or when the token must be removed entirely. The entropy term maintains spatial focus and enhances video quality. The objective is optimized over latents during inference, while the model weights remain fixed.